# Comparative Explanations: Explanation Guided Decision Making for Human-in-the-Loop Preference Selection

### v1

@@ 3.1 Approach
7:         $\mathcal{X}_{\text{lhs}}\leftarrow\{\mathbf{x}_{1}^{\prime},\ldots,\mathbf{x}_{N}^{\prime}\}\sim\text{LHS}(\mathbf{x},\text{radius}=r)$triangleright Generate explanation set
8:         $\mu_{f}(\mathbf{x}_{s}),\sigma_{f}(\mathbf{x}_{s})\leftarrow\mathcal{M}(\mathbf{x}_{s}),\quad\forall\mathbf{x}_{s}\in\mathcal{X}_{\text{lhs}}$triangleright GPf model evaluation
9:         $\mu_{g}(\mu_{f}(\mathbf{x}_{s})),\sigma_{g}(\mu_{f}(\mathbf{x}_{s}))\leftarrow\mathcal{M}_{pref}(\mathcal{M}(\mathbf{x}_{s})),\quad\forall\mathbf{x}_{s}\in\mathcal{X}_{\text{lhs}}$triangleright GPg model evaluation
10:         $\boldsymbol{\phi_{x}}\leftarrow\text{InputFeatureImportance}(\mathcal{M},\mathcal{X}_{\text{lhs}})$triangleright Input feature importance
11:         $\boldsymbol{\phi_{y}}\leftarrow\text{OutcomeImportance}(\mathcal{M}_{pref},\mu_{f}(\mathbf{x}_{s}))$triangleright Outcome importance
@@
13:     $\mathcal{H}\leftarrow\{\text{High}(\boldsymbol{\phi_{x}}),\text{High}(\boldsymbol{\phi_{y}})\}$triangleright Why explanations
14:     $\mathcal{L}\leftarrow\{\text{Low}(\boldsymbol{\phi_{x}}),\text{Low}(\boldsymbol{\phi_{y}})\}$triangleright Why not explanations
15:     $\text{CompMatrix}\leftarrow\text{Combine}(\mathcal{H}_{WhyExp},\mathcal{L}_{WhyNotExp})$triangleright Comparative matrix
@@ 3.2 Comparative Explanations
Mathematically, this approach can be described as follows: LHS generates $N$ points $\mathbf{u}_{i}$ uniformly within a unit cube $[0,1]^{d}$. These points are then transformed into a spherical region by first normalizing them to lie on the surface of a sphere and then scaling them by a radius $r\sim U(0,1)^{1/d}$, ensuring uniform sampling within the volume of the sphere. Finally, the points are translated to be centered at $\mathbf{x}$: $\mathbf{x}_{i}^{\prime}=\mathbf{x}+r\frac{\mathbf{u}_{i}}{\|\mathbf{u}_{i}\|}$, where $\mathbf{u}_{i}$ is the LHS-sampled point, $r$ is the radius, and $\|\mathbf{u}_{i}\|$ is the Euclidean norm of $\mathbf{u}_{i}$.
@@ 4.1 Optimization Benchmark Functions/Dataset
The DTLZ2 problem is a $d$-dimensional function with 5 decision variables (inputs) and 4 outputs, evaluated over the domain $[0,1]^{d}$ [9]. We adapt this problem for PBO by modifying its objective to maximize the sum of the first three outputs, reflecting decision-making scenarios where certain dimensions hold greater importance.
@@
The ZDT1 problem involves 5 decision variables and 2 objectives, evaluated over the domain $[0,1]^{d}$. This function tests the algorithm’s ability to handle trade-offs between a linearly scaled objective and a non-linear function that depends on the remaining variables [38]. We adapt ZDT1 to the PBO framework by modifying the objective to maximize the sum of its two outputs.
@@ 5.3 Evaluation with Humans
A Shapiro-Wilk test confirmed that the distributions of mean utility were non-normal for all groups: ideal agent ($W=0.9,p<0.01$), users without explanations ($W=0.8,p<0.01$), and users with explanations ($W=0.8,p<0.01$). The results (Fig. 7) indicate that users with explanations performed on par with the ideal selection agent, achieving a mean utility of $M=1.83$ (ideal agent: ±$\pm 0.075$, users with explanations: ±$\pm 0.070$). In contrast, users without explanations had a lower mean utility of $M=1.69\pm 0.036$, suggesting that the absence of explanations negatively impacted decision quality. A Kruskal-Wallis test confirmed that this difference was statistically significant ($X^{2}(2,N=5)=18.87,p<0.01$).

### v2

@@ 3.1 Approach
7:   $\mathcal{X}_{\text{lhs}}\leftarrow\{\mathbf{x}_{1}^{\prime},\ldots,\mathbf{x}_{N}^{\prime}\}\sim\text{LHS}(\mathbf{x},\text{radius}=r)$ $\triangleright$ Generate explanation set
8:   $\mu_{f}(\mathbf{x}_{s}),\sigma_{f}(\mathbf{x}_{s})\leftarrow\mathcal{M}(\mathbf{x}_{s}),\quad\forall\mathbf{x}_{s}\in\mathcal{X}_{\text{lhs}}$ $\triangleright$ GPf model evaluation
9:   $\mu_{g}(\mu_{f}(\mathbf{x}_{s})),\sigma_{g}(\mu_{f}(\mathbf{x}_{s}))\leftarrow\mathcal{M}_{pref}(\mathcal{M}(\mathbf{x}_{s})),\quad\forall\mathbf{x}_{s}\in\mathcal{X}_{\text{lhs}}$ $\triangleright$ GPg model evaluation
10:   $\boldsymbol{\phi_{x}}\leftarrow\text{InputFeatureImportance}(\mathcal{M},\mathcal{X}_{\text{lhs}})$ $\triangleright$ Input feature importance
11:   $\boldsymbol{\phi_{y}}\leftarrow\text{OutcomeImportance}(\mathcal{M}_{pref},\mu_{f}(\mathbf{x}_{s}))$ $\triangleright$ Outcome importance
@@
13:  $\mathcal{H}\leftarrow\{\text{High}(\boldsymbol{\phi_{x}}),\text{High}(\boldsymbol{\phi_{y}})\}$ $\triangleright$ Why explanations
14:  $\mathcal{L}\leftarrow\{\text{Low}(\boldsymbol{\phi_{x}}),\text{Low}(\boldsymbol{\phi_{y}})\}$ $\triangleright$ Why not explanations
15:  $\text{CompMatrix}\leftarrow\text{Combine}(\mathcal{H}_{WhyExp},\mathcal{L}_{WhyNotExp})$ $\triangleright$ Comparative matrix
@@ 3.2 Comparative Explanations
Mathematically, this approach can be described as follows: LHS generates $N$ points $\mathbf{u}_{i}$ uniformly within a unit cube $0,1]^{d}$. These points are then transformed into a spherical region by first normalizing them to lie on the surface of a sphere and then scaling them by a radius $r\sim U(0,1)^{1/d}$, ensuring uniform sampling within the volume of the sphere. Finally, the points are translated to be centered at $\mathbf{x}$: $\mathbf{x}_{i}^{\prime}=\mathbf{x}+r\frac{\mathbf{u}_{i}}{\|\mathbf{u}_{i}\|}$, where $\mathbf{u}_{i}$ is the LHS-sampled point, $r$ is the radius, and $\|\mathbf{u}_{i}\|$ is the Euclidean norm of $\mathbf{u}_{i}$.
@@ 4.1 Optimization Benchmark Functions/Dataset
The DTLZ2 problem is a $d$-dimensional function with 5 decision variables (inputs) and 4 outputs, evaluated over the domain $0,1]^{d} [$9]. We adapt this problem for PBO by modifying its objective to maximize the sum of the first three outputs, reflecting decision-making scenarios where certain dimensions hold greater importance.
@@
The ZDT1 problem involves 5 decision variables and 2 objectives, evaluated over the domain $0,1]^{d}$. This function tests the algorithm’s ability to handle trade-offs between a linearly scaled objective and a non-linear function that depends on the remaining variables [38]. We adapt ZDT1 to the PBO framework by modifying the objective to maximize the sum of its two outputs.
@@ 5.3 Evaluation with Humans
A Shapiro-Wilk test confirmed that the distributions of mean utility were non-normal for all groups: ideal agent ($W=0.9,p<0.01$), users without explanations ($W=0.8,p<0.01$), and users with explanations ($W=0.8,p<0.01$). The results (Fig. 7) indicate that users with explanations performed on par with the ideal selection agent, achieving a mean utility of $M=1.83$ (ideal agent: $\pm 0.075$, users with explanations: $\pm 0.070$). In contrast, users without explanations had a lower mean utility of $M=1.69\pm 0.036$, suggesting that the absence of explanations negatively impacted decision quality. A Kruskal-Wallis test confirmed that this difference was statistically significant ($X^{2}(2,N=5)=18.87,p<0.01$).